# Towards Agentic Intelligence for Materials Science

### v1

@@ 3.3.3 Synthesis Route Generation
Synthesis route generation refers to constructing feasible pathways to synthesize a target material. In materials science, it covers various aspects, such as choosing suitable precursors, suggesting reaction paths, and making predictions about key reaction conditions like time, temperature, and reaction environment 419. In practice, synthesis is often the dominant bottleneck: feasibility and process constraints are what separate generated candidates from experimentally validated materials. Graph-based methodologies were first applied to depict solid-state processes and employed pathfinding to ascertain viable synthesis pathways 239. Researchers have also used machine learning to directly predict experimental conditions from data in the literature. This has shown links between the stability of precursors and the temperatures or periods of synthesis156. In addition to reactions, models can suggest sets of precursors based on how similar the materials are, which makes recipes for inorganic synthesis more accurate147. At the system level, autonomous laboratories are already combining AI planning with robotics and active learning. For example, A-Lab can design, run, and refine synthesis strategies for new inorganic compounds iteratively again342. What makes such platforms non-reactive is closed-loop execution, where synthesis outcomes update models and trigger replanning, rather than ending at a one-shot proposal.
Synthesis planning is a complex task requiring the collaborative optimization of advanced algorithms such as deep learning, reinforcement learning, and uncertainty-aware search 408. The aforementioned approaches reveal that synthesis planning is evolving toward a data-driven closed-loop system. The systems must not only generate synthesis paths tailored to specific requirements but also perform real-time updates based on synthesis outcomes 328. These loops can supply discovery-level feedback signals that support end-to-end credit assignment across upstream components.
@@ 3.4 Optimization and Verification
3.4.1 Material Discovery Process Optimization
As an instance of the optimization lab, a self-driving platform, Ada 233, has been developed; an AI-driven lab for the synthesis of palladium thin films that uses a Bayesian optimization algorithm to autonomously optimize the process in a closed loop where objectives are measured on experimental outcomes. By mapping the conductivity-temperature trade-off, the AI model identifies the optimal low-temperature, high-conductivity formulations, which were later successfully validated in a scalable manufacturing process. Later,  Chitturi et al. 74 demonstrated an autonomous material search platform by developing a multi property version of Bayesian algorithm execution (BAX), which transforms the user-defined targets into acquisition functions to guide the autonomous targeted material discovery. Reported evaluations emphasize efficiency in locating target materials; for discovery-oriented settings, these should be interpreted through outcome-level metrics (e.g., experimental sample efficiency), not benchmark-style scores alone. Benchmarks showed that variants of BAX outperform the strong baseline methods in efficiently locating target materials, but only in low-dimensional settings 381.
SPINBOT 412, an autonomous platform employing Bayesian optimization, optimizes perovskite film fabrication based on photoluminescence feedback. Similarly, ARES 64 utilized Bayesian optimization to accelerate carbon nanotube synthesis, achieving an 8-fold increase in the growth rate of carbon nanotubes, demonstrating the effectiveness of closed-loop optimization.
@@ 3.4.2 Simulation and AI-based Verification
Fare et al. 111 proposed a method by leveraging the multi-output Gaussian process that dynamically fuses information from sources of varying accuracy and cost (e.g., simulations and experiments). The optimization task is driven by Targeted Variance Reduction with Expected Improvement (TVR-EI), a multi-fidelity Bayesian algorithm, verified by performing computational simulations. More broadly, verification in materials discovery is inherently multi-fidelity (approximate vs accurate), so pipeline-centric systems must allocate verification budgets adaptively. Merchant et al. 242 demonstrated Graph Networks for Materials Exploration (GNOME) for large-scale materials discovery by combining GNNs with Density Functional Theory (DFT) simulations.
For the autonomous investigation of the mechanism of catalytic reactions, Lan et al. 203 applied deep reinforcement learning methods by developing a high-throughput framework coupled with first-principles DFT simulations. More recently, Wen et al. 387 showcased the integration of Cartesian Atomic Moment Potential (CAMP), a machine-learned interatomic potential (MLIP) built entirely in Cartesian space, in closed-loop simulations, as the next generation MLIP primitive. They have demonstrated that CAMP offers systematic improvements through utilizing physically motivated atomic moment tensors and their tensor products within a graph neural network framework to capture local atomic environments and higher-order interactions. From a pipeline-centric view, improved MLIPs matter because they reduce verification cost and expand the feasible horizon of simulation-driven loops, making longer-horizon, agentic planning more practical.
@@ 3.4.3 Agent-based Closed-loop Labs
Closed-loop labs accelerate research by leveraging artificial intelligence through the integration of AI with automated experiments 196, 5, 315. The “mobile robotic chemist” was presented by Burger et al. 57, employing a free-roaming robot, as an autonomous agent, to automate the researcher’s tasks. The robot was tasked with searching for an improved photocatalyst, driven by a batched Bayesian search algorithm exploring a ten-variable experimental space, and Gaussian process regression was used to build a predictive model. The “Artificial Chemist”, a self-driven agent, was developed by Epps et al. 104 by coupling an autonomous microfluidic flow reactor with machine learning to optimize quantum dot synthesis. Later Sadeghi et al. 306 developed a self-driving fluidic lab by integrating microfluidics, automation, and Bayesian optimization for the autonomous synthesis of lead-free perovskite nanocrystals. Kusne et al. 200 presented CAMEO, closed-loop autonomous materials exploration and optimization, an autonomous agent designed to discover new inorganic materials, driven by Bayesian active learning to explore the complex relationship among the composition, structural phase, and properties of the material. The “A-Lab” was developed by Szymanski et al. 342, an autonomous platform for the synthesis of novel solid-state inorganic materials. The A-Lab platform employed natural language processing models trained on literature to propose initial recipes, and optimization was carried out by ARROWS 341, an active learning algorithm that learns from previous experimental outcomes. Upon failure of synthesis, (ARROWS) uses ab initio computed reaction energies to predict and optimize alternative reaction pathways. More recently, Omidvar et al. 265 presented an agent-based closed-loop platform for the synthesis of perovskite solid solutions by combining ML-guided composition prediction, automated solid-state synthesis, and high-throughput characterization.
@@
LLMs have broadened self-directed material exploration. In the realm of the application of LLMs in chemical synthesis, Coscientist, an autonomous chemical research framework, has been developed by employing GPT-4-based agents to orchestrate chemical experiments 46. Additionally, in materials characterization, LLMs are currently employed not only for experimental design but also for instrument control and automation via code generation. For example, one implementation demonstrated that an LLM could independently write executable scripts (e.g., Python and LabVIEW) to control laboratory instruments from natural language prompts, showcasing genuine hardware control with minimal human involvement 223. In recent times, LLMs have been used to achieve full laboratory autonomy by bringing together experiment design, instrument operation, data collection, and analysis into one system. More recently, LLM-centered systems have been explored for tighter end-to-end loops that combine experiment design, instrument operation, data collection, and analysis. Xie et al. 395 described a comprehensive system in which the LLM not only produced control code but also analyzed experimental outcomes and enhanced subsequent measurement circumstances, thereby establishing a closed-loop connection between thinking and execution. These research efforts clearly show that we are moving toward experimental autonomy based on natural language processing. In this new way, LLMs are not just instruments for communication; they are also agents for coding, control, and reasoning that connect human purpose with machine execution.
@@
To integrate human experts with machine learning, Adams et al. 8 developed a human-in-the-loop Bayesian framework where expert feedback is incorporated as priors to guide automated phase mapping experiments. BOARS addressed the rigidity of predefined targets in autonomous experiments by developing a human-in-the-loop workflow 43. The optimization strategy allows a human operator to dynamically define and refine the scientific objective during the experiment by upvoting or downvoting real-time spectral measurements, successfully locating the human-desired optimum in significantly less time compared to manual optimization.
@@ 3.5 Data and Knowledge
One of the long-standing challenges in applying AI to material science is the scarcity and heterogeneity of material data and knowledge. This becomes more challenging for an end-to-end agentic LLM for materials science, especially considering the costs associated with evaluating the outputs in the real-world, never-ending experimental environment.
@@ 3.5.1 Data scarcity and heterogeneity
The advancement of data-driven materials science is characterized by the collaborative endeavors of all stakeholders to address the challenges of data scarcity and data heterogeneity 399, 241. To solve this problem, systems like MatMiner 383 and Robocrystallographer 117 give us the tools we need to turn different types of materials into standardized numerical descriptors that ML can use. MatMiner has a full library of features that cover a wide range of attributes. Robocrystallographer, on the other hand, is made to create interpretable features from crystallographic data.
OPTIMADE 20 provides a unified API enabling researchers to query diverse material databases simultaneously, thereby partially addressing data heterogeneity issues. Ghiringhelli et al. 125 proposes a hierarchical metadata approach,FAIR, which stands for Findable, Accessible, Interoperable, and Reusable, advocating for a unified and streamlined philosophy for computational and experimental data management.
The LeMatTraj project 295 directly solves the problem of not having enough data by combining nearly 120 million high-quality atomic configurations from the primary repository into a single dataset. This is in addition to the infrastructure standards. This huge collection has the high-quality, standardized data that is needed to develop strong machine learning interatomic potentials.
@@
Data augmentation techniques are crucial for mitigating data scarcity and heterogeneity in materials modeling, evolving from basic perturbation strategies to complex physics-informed generative methods. Early studies, such as OC20 65, supplemented DFT datasets through random structural perturbations, short-timescale molecular dynamics simulations, and electronic structure analysis. It then slowly changed into physics-informed augmentation strategies: Szymanski et al. 340 included experimental errors (lattice strain, crystal texture) in the XRD pattern generation process, thus linking simulations to experiments. Gibson et al. 127 proposed a physically motivated, computationally cheap perturbation technique that augments training data to improve the accuracy of unrelaxed structure prediction by 66% and directly addressing the domain mismatch problem in crystal structure prediction. Based on this idea, Dinic et al. 97 presented an ML model capable of predicting the crystal energy response to global strain by using available elasticity data to augment the dataset, which greatly improved energy forecasts for structures that were not perfectly straight. In addition to physics-motivated methods, generative techniques that produce high-quality in-silico data to augment the original dataset have developed.Chung et al. 81 proposed a GAN-based data synthesis method to produce spectral samples to address the class imbalance problem, thereby enhancing the performance of the ML model trained on the augmented dataset. MatWheel 214 introduced a framework that trains the material property prediction model using the synthetic data generated by the conditional generative model, showing its potential in addressing extreme data-scarce scenarios in materials science.
@@
Due to the trade-off between cost and accuracy, materials science involves measurements and simulation data with varying degrees of precision. To enable the fusion learning of existing data with different levels of accuracy, researchers have proposed multi-fidelity learning. Its objective is to maximize the utilization of existing data for high-accuracy model predictions while minimizing additional computational and human resource costs. Pilania et al. 279 introduced this approach using co-kriging to fuse low-cost, low-accuracy DFT calculations with sparse high-fidelity data, achieving accurate electronic band gap predictions at significantly reduced computational cost. Further, MODNet 92 replaced Gaussian process regression with DNN while learning about the differences between high- and low-quality values, which shows significant improvement in the results compared to learning on the sole high-quality experimental data. In contrast, MD-HIT 212 addresses the redundancy problem inherent in multi-source datasets, employing greedy incremental algorithms to filter overlapping samples and prevent overestimation of performance, thereby enabling more realistic evaluation and improved generalization to out-of-distribution materials.
@@
Few-shot learning, which is training a model on a small number of labeled examples, can help models quickly learn new tasks, especially when there is not a lot of good training data available. Meta-MGNN 143 was the first works creating meta-learning frameworks for predicting molecular characteristics. This framework allows models to quickly adapt to new attributes with training on a few instances, and also use self-supervised learning from molecular structures to make up for the lack of labeled data. AttFPGNN-MAML 288 built on this by combining hybrid molecular representations fingerprints and graph neural networks with ProtoMAML meta-learning. This made it easier to adapt to new prediction tasks. Most recently, MolFeSCue 415 moved the field forward by adding dynamic contrastive loss functions that let pretrained models learn useful representations from very unbalanced and small amounts of training data. It addresses the challenge of data scarcity in drug discovery and materials research.
@@ 3.5.2 Knowledge integration
Physics-informed neural networks (PINNs) are a type of general-purpose function approximator that can embed knowledge of any physical laws governing a given dataset into the learning process. Chen et al. 71 employs PINNs for the solution of inverse scattering problems that retrieves the effective permittivity parameters of a number of finite-size scattering systems in photonic metamaterials and nanooptics. Fang and Zhan 110 applied deep physics-informed neural networks for electromagnetic metamaterial design, which can be utilized in dealing with various practical problems such as cloaking, rotators, concentrators, etc. Zhang et al. 411 employ PINNs for defect characterization in mechanical materials, using equilibrium PDEs and constitutive laws to identify internal voids/inclusions from boundary displacement measurements.
@@
ElemNet 162 and CrabNet 368 only use the elemental composition of the material as input. They automatically learn chemical relationships and composition-property correlations between elements based on chemical relationships in the periodic table, without requiring any manually designed descriptors. In addition, GNNs such as MEGNet 69, ALIGNN 78, ROOST 130, and GnoME 242 model molecules and crystals by treating atoms as nodes, bonds as edges, and selectively adding global features such as temperature or pressure. This approach allows deep learning algorithms to make very accurate predictions of features such as formation energy, band gap, and elastic modulus. Based on this general modeling, many new studies have incorporated additional prior knowledge of material structure features by introducing new encoding modules or learning objectives to integrate structural knowledge: M3GNet 68 introduced a many-body computation module to calculate the three-body and many-body interactions; NequIP 33 proposed E(3)-equivariant neural networks using geometric tensors and equivariant convolutions based on tensor products with spherical harmonics, which introduces feature encoding of the rotational and translational symmetries of the material structure. Allegro 250 removes atom-centered message forwarding while maintaining equivariance by employing learned representations in iterated tensor products.
@@
As discussed in Section 3.2, extensive knowledge bases and knowledge graphs have been constructed from various sources, including academic articles and experimental analyses. This organized knowledge in databases, covering multiple kinds of knowledge in materials science, can then be used to enhance model performance in material prediction tasks. Zhao et al. 424 is specifically designed to address the challenges of data heterogeneity in materials science. It uses knowledge graphs as a semantic integration center and combines ontology mapping and automatic annotation to extract information from academic articles. Both KCL 108, KANO 109, ESNet 152 construct an element-centric knowledge graph capturing intrinsic physical and chemical attributes of elements and leverages it in a multimodal or contrastive learning framework to unify structural features with semantic chemical knowledge for improved molecular and materials property prediction. Recent innovations combine LLM with knowledge integration: MOFs-KG 19 createed the first comprehensive KGQA benchmark specifically for materials science (MOFs) with 644 complex natural language questions, demonstrating ChatGPT’s systematic application to domain-specific knowledge graphs.
@@
Pretrained models have been demonstrated to be effective in natural language processing, especially in data scarcity scenarios. Through training with limited data, pretrained models can be transferred to a specific task while showing strong performance. In materials science, pretrained model are also widely used. Early work, like Mat2Vec 358, learned word embeddings in material literature without human supervision, showing a surprising capability to identify semantics in material science text. This led to the development of material-specific language models such as MatSciBERT 145 and MaterialBERT 318. They collect a large corpus of material science for continual training without changing the BERT architecture. Compared to vanilla BERT, they show superior performance in identifying material-related property data and phrases in the literature. This conclusion contributes to the emergence of various BERT variations that were particular to certain materials subfields, such as BatteryBERT 155 for battery-related material, OpticalBERT 423 for optical materials, PolyBERT 195 for polymers materials, and SteelBERT 352 for metallurgy materials. UniMat 261 propelled the field forward by merging multiple modes (atomic structure, XRD patterns, and composition) into strong joint embeddings. This made it easier to find novel materials. SmileyLlama 61 is further eolved with Direct Preference Optimization(DPO) to optimize the structure generation. Recent ideas deal with structural limitations: Huang et al. 153 proposed a self-supervised and multimodal pretraining technique that can predict performance based only on stoichiometry, without explicit understanding of the atomic structure. MELT188 was the first to employ curriculum-based continuous pretraining, which used semantic knowledge graphs to convert models from broad material concepts to specific material concepts in a systematic way.

### v2

@@ 3.3.3 Synthesis Route Generation
Synthesis route generation refers to constructing feasible pathways to synthesize a target material. In materials science, it covers various aspects, such as choosing suitable precursors, suggesting reaction paths, and making predictions about key reaction conditions like time, temperature, and reaction environment 422. In practice, synthesis is often the dominant bottleneck: feasibility and process constraints are what separate generated candidates from experimentally validated materials. Graph-based methodologies were first applied to depict solid-state processes and employed pathfinding to ascertain viable synthesis pathways 241. Researchers have also used machine learning to directly predict experimental conditions from data in the literature. This has shown links between the stability of precursors and the temperatures or periods of synthesis157. In addition to reactions, models can suggest sets of precursors based on how similar the materials are, which makes recipes for inorganic synthesis more accurate148. At the system level, autonomous laboratories are already combining AI planning with robotics and active learning. For example, A-Lab can design, run, and refine synthesis strategies for new inorganic compounds iteratively again344. What makes such platforms non-reactive is closed-loop execution, where synthesis outcomes update models and trigger replanning, rather than ending at a one-shot proposal.
Synthesis planning is a complex task requiring the collaborative optimization of advanced algorithms such as deep learning, RL, and uncertainty-aware search 411. The aforementioned approaches reveal that synthesis planning is evolving toward a data-driven closed-loop system. The systems must not only generate synthesis paths tailored to specific requirements but also perform real-time updates based on synthesis outcomes 330. These loops can supply discovery-level feedback signals that support end-to-end credit assignment across upstream components.
@@ 3.4 Optimization and Verification
3.4.1 Materials Discovery Process Optimization
As an instance of the optimization lab, a self-driving platform, Ada 235, has been developed; an AI-driven lab for the synthesis of palladium thin films that uses a Bayesian optimization algorithm to autonomously optimize the process in a closed loop where objectives are measured on experimental outcomes. By mapping the conductivity-temperature trade-off, the AI model identifies the optimal low-temperature, high-conductivity formulations, which were later successfully validated in a scalable manufacturing process. Later,  Chitturi et al. 75 demonstrated an autonomous material search platform by developing a multi-property version of Bayesian algorithm execution (BAX), which transforms the user-defined targets into acquisition functions to guide the autonomous targeted materials discovery. Reported evaluations emphasize efficiency in locating target materials; for discovery-oriented settings, these should be interpreted through outcome-level metrics (e.g., experimental sample efficiency), not benchmark-style scores alone. Benchmarks showed that variants of BAX outperform the strong baseline methods in efficiently locating target materials, but only in low-dimensional settings 384.
SPINBOT 415, an autonomous platform employing Bayesian optimization, optimizes perovskite film fabrication based on photoluminescence feedback. Similarly, Chang et al. 65 developed an Autonomous Research System (ARES) utilized Bayesian optimization to accelerate carbon nanotube synthesis, achieving an 8-fold increase in the growth rate of carbon nanotubes, demonstrating the effectiveness of closed-loop optimization. Besides Bayesian optimization, RLHF is also used in M4olGen 217 to achieve multi-objective optimization on property-constrained molecular generation.
@@ 3.4.2 Simulation and AI-based Verification
Fare et al. 112 proposed a method by leveraging the multi-output Gaussian process that dynamically fuses information from sources of varying accuracy and cost (e.g., simulations and experiments). The optimization task is driven by Targeted Variance Reduction with Expected Improvement (TVR-EI), a multi-fidelity Bayesian algorithm, verified by performing computational simulations. More broadly, verification in materials discovery is inherently multi-fidelity (approximate vs accurate), so pipeline-centric systems must allocate verification budgets adaptively. Merchant et al. 244 demonstrated Graph Networks for Materials Exploration (GNOME) for large-scale materials discovery by combining GNNs with Density Functional Theory (DFT) simulations.
For the autonomous investigation of the mechanism of catalytic reactions, Lan et al. 204 applied deep RL methods by developing a high-throughput framework coupled with first-principles DFT simulations. More recently, Wen et al. 390 showcased the integration of Cartesian Atomic Moment Potential (CAMP), a machine-learned interatomic potential (MLIP) built entirely in Cartesian space, in closed-loop simulations, as the next generation MLIP primitive. They have demonstrated that CAMP offers systematic improvements through utilizing physically motivated atomic moment tensors and their tensor products within a graph neural network framework to capture local atomic environments and higher-order interactions. From a pipeline-centric view, improved MLIPs matter because they reduce verification cost and expand the feasible horizon of simulation-driven loops, making longer-horizon, agentic planning more practical.
@@ 3.4.3 Agent-based Closed-loop Labs
Closed-loop labs accelerate research through the integration of AI with automated and self-driving experiments 197, 5, 317, 22, 356. The “mobile robotic chemist” was presented by Burger et al. 58, employing a free-roaming robot, as an autonomous agent, to automate the researcher’s tasks. The robot was tasked with searching for an improved photocatalyst, driven by a batched Bayesian search algorithm exploring a ten-variable experimental space, and Gaussian process regression was used to build a predictive model. The “Artificial Chemist”, a self-driven agent, was developed by Epps et al. 105 by coupling an autonomous microfluidic flow reactor with machine learning to optimize quantum dot synthesis. Later Sadeghi et al. 308 developed a self-driving fluidic lab by integrating microfluidics, automation, and Bayesian optimization for the autonomous synthesis of lead-free perovskite nanocrystals. Kusne et al. 201 presented CAMEO, closed-loop autonomous materials exploration and optimization, an autonomous agent designed to discover new inorganic materials, driven by Bayesian active learning to explore the complex relationship among the composition, structural phase, and properties of the material. The “A-Lab” was developed by Szymanski et al. 344, an autonomous platform for the synthesis of novel solid-state inorganic materials. The A-Lab platform employed natural language processing models trained on literature to propose initial recipes, and optimization was carried out by ARROWS 343, an active learning algorithm that learns from previous experimental outcomes. Upon failure of synthesis, (ARROWS) uses ab initio computed reaction energies to predict and optimize alternative reaction pathways. More recently, Omidvar et al. 267 presented an agent-based closed-loop platform for the synthesis of perovskite solid solutions by combining ML-guided composition prediction, automated solid-state synthesis, and high-throughput characterization.
@@
LLMs have broadened self-directed materials exploration. In the realm of the application of LLMs in chemical synthesis, Coscientist, an autonomous chemical research framework, has been developed by employing GPT-4-based agents to orchestrate chemical experiments 47. Additionally, in materials characterization, LLMs are currently employed not only for experimental design but also for instrument control and automation via code generation. For example, one implementation demonstrated that an LLM could independently write executable scripts (e.g., Python and LabVIEW) to control laboratory instruments from natural language prompts, showcasing genuine hardware control with minimal human involvement 225. In recent times, LLMs have been used to achieve full laboratory autonomy by bringing together experiment design, instrument operation, data collection, and analysis into one system. More recently, LLM-centered systems have been explored for tighter end-to-end loops that combine experiment design, instrument operation, data collection, and analysis. Xie et al. 398 described a comprehensive system in which the LLM not only produced control code but also analyzed experimental outcomes and enhanced subsequent measurement circumstances, thereby establishing a closed-loop connection between thinking and execution. These research efforts clearly show that we are moving toward experimental autonomy based on natural language processing. In this new way, LLMs are not just instruments for communication; they are also agents for coding, control, and reasoning that connect human purpose with machine execution.
@@
To integrate human experts with machine learning, Adams et al. 8 developed a human-in-the-loop Bayesian framework where expert feedback is incorporated as priors to guide automated phase mapping experiments. BOARS addressed the rigidity of predefined targets in autonomous experiments by developing a human-in-the-loop workflow 44. The optimization strategy allows a human operator to dynamically define and refine the scientific objective during the experiment by upvoting or downvoting real-time spectral measurements, successfully locating the human-desired optimum in significantly less time compared to manual optimization.
@@ 3.5 Data and Knowledge
One of the long-standing challenges in applying AI to materials science is the scarcity and heterogeneity of material data and knowledge. This becomes more challenging for an end-to-end agentic LLM for materials science, especially considering the costs associated with evaluating the outputs in the real-world, never-ending experimental environment.
@@ 3.5.1 Data scarcity and heterogeneity
The advancement of data-driven materials science is characterized by the collaborative endeavors of all stakeholders to address the challenges of data scarcity and data heterogeneity 402, 243. To solve this problem, systems like MatMiner 386 and Robocrystallographer 118 give us the tools we need to turn different types of materials into standardized numerical descriptors that ML can use. MatMiner has a full library of features that cover a wide range of attributes. Robocrystallographer, on the other hand, is made to create interpretable features from crystallographic data.
OPTIMADE 20 provides a unified API enabling researchers to query diverse materials databases simultaneously, thereby partially addressing data heterogeneity issues. Ghiringhelli et al. 126 proposes a hierarchical metadata approach,FAIR, which stands for Findable, Accessible, Interoperable, and Reusable, advocating for a unified and streamlined philosophy for computational and experimental data management.
The LeMatTraj project 297 directly solves the problem of not having enough data by combining nearly 120 million high-quality atomic configurations from the primary repository into a single dataset. This is in addition to the infrastructure standards. This huge collection has the high-quality, standardized data that is needed to develop strong machine learning interatomic potentials.
@@
Data augmentation techniques are crucial for mitigating data scarcity and heterogeneity in materials modeling, evolving from basic perturbation strategies to complex physics-informed generative methods. Early studies, such as OC20 66, supplemented DFT datasets through random structural perturbations, short-timescale molecular dynamics simulations, and electronic structure analysis. It then slowly changed into physics-informed augmentation strategies: Szymanski et al. 342 included experimental errors (lattice strain, crystal texture) in the XRD pattern generation process, thus linking simulations to experiments. Gibson et al. 128 proposed a physically motivated, computationally cheap perturbation technique that augments training data to improve the accuracy of unrelaxed structure prediction by 66% and directly addressing the domain mismatch problem in crystal structure prediction. Based on this idea, Dinic et al. 98 presented an ML model capable of predicting the crystal energy response to global strain by using available elasticity data to augment the dataset, which greatly improved energy forecasts for structures that were not perfectly straight. In addition to physics-motivated methods, generative techniques that produce high-quality in-silico data to augment the original dataset have developed.Chung et al. 82 proposed a GAN-based data synthesis method to produce spectral samples to address the class imbalance problem, thereby enhancing the performance of the ML model trained on the augmented dataset. MatWheel 215 introduced a framework that trains the materials property prediction model using the synthetic data generated by the conditional generative model, showing its potential in addressing extreme data-scarce scenarios in materials science.
@@
Due to the trade-off between cost and accuracy, materials science involves measurements and simulation data with varying degrees of precision. To enable the fusion learning of existing data with different levels of accuracy, researchers have proposed multi-fidelity learning. Its objective is to maximize the utilization of existing data for high-accuracy model predictions while minimizing additional computational and human resource costs. Pilania et al. 281 introduced this approach using co-kriging to fuse low-cost, low-accuracy DFT calculations with sparse high-fidelity data, achieving accurate electronic band gap predictions at significantly reduced computational cost. Further, MODNet 93 replaced Gaussian process regression with DNN while learning about the differences between high- and low-quality values, which shows significant improvement in the results compared to learning on the sole high-quality experimental data. In contrast, MD-HIT 213 addresses the redundancy problem inherent in multi-source datasets, employing greedy incremental algorithms to filter overlapping samples and prevent overestimation of performance, thereby enabling more realistic evaluation and improved generalization to out-of-distribution materials.
@@
Few-shot learning, which is training a model on a small number of labeled examples, can help models quickly learn new tasks, especially when there is not a lot of good training data available. Meta-MGNN 144 was the first works creating meta-learning frameworks for predicting molecular characteristics. This framework allows models to quickly adapt to new attributes with training on a few instances, and also use self-supervised learning from molecular structures to make up for the lack of labeled data. AttFPGNN-MAML 290 built on this by combining hybrid molecular representations fingerprints and graph neural networks with ProtoMAML meta-learning. This made it easier to adapt to new prediction tasks. Most recently, MolFeSCue 418 moved the field forward by adding dynamic contrastive loss functions that let pre-trained models learn useful representations from very unbalanced and small amounts of training data. It addresses the challenge of data scarcity in drug discovery and materials research.
@@ 3.5.2 Knowledge integration
Physics-informed neural networks (PINNs) are a type of general-purpose function approximator that can embed knowledge of any physical laws governing a given dataset into the learning process. Chen et al. 72 employs PINNs for the solution of inverse scattering problems that retrieves the effective permittivity parameters of a number of finite-size scattering systems in photonic metamaterials and nanooptics. Fang and Zhan 111 applied deep physics-informed neural networks for electromagnetic metamaterial design, which can be utilized in dealing with various practical problems such as cloaking, rotators, concentrators, etc. Zhang et al. 414 employ PINNs for defect characterization in mechanical materials, using equilibrium PDEs and constitutive laws to identify internal voids/inclusions from boundary displacement measurements.
@@
ElemNet 163 and CrabNet 371 only use the elemental composition of the material as input. They automatically learn chemical relationships and composition-property correlations between elements based on chemical relationships in the periodic table, without requiring any manually designed descriptors. In addition, GNNs such as MEGNet 70, ALIGNN 79, ROOST 131, and GnoME 244 model molecules and crystals by treating atoms as nodes, bonds as edges, and selectively adding global features such as temperature or pressure. This approach allows deep learning algorithms to make very accurate predictions of features such as formation energy, band gap, and elastic modulus. Based on this general modeling, many new studies have incorporated additional prior knowledge of material structure features by introducing new encoding modules or learning objectives to integrate structural knowledge: M3GNet 69 introduced a many-body computation module to calculate the three-body and many-body interactions; NequIP 34 proposed E(3)-equivariant neural networks using geometric tensors and equivariant convolutions based on tensor products with spherical harmonics, which introduces feature encoding of the rotational and translational symmetries of the material structure. Allegro 252 removes atom-centered message forwarding while maintaining equivariance by employing learned representations in iterated tensor products.
@@
As discussed in Section 3.2, extensive knowledge bases and knowledge graphs have been constructed from various sources, including academic articles and experimental analyses. This organized knowledge in databases, covering multiple kinds of knowledge in materials science, can then be used to enhance model performance in material property prediction tasks. Zhao et al. 427 is specifically designed to address the challenges of data heterogeneity in materials science. It uses knowledge graphs as a semantic integration center and combines ontology mapping and automatic annotation to extract information from academic articles. Both KCL 109, KANO 110, ESNet 153 construct an element-centric knowledge graph capturing intrinsic physical and chemical attributes of elements and leverages it in a multimodal or contrastive learning framework to unify structural features with semantic chemical knowledge for improved molecular and materials property prediction. Recent innovations combine LLM with knowledge integration: MOFs-KG 19 createed the first comprehensive KGQA benchmark specifically for materials science (MOFs) with 644 complex natural language questions, demonstrating ChatGPT’s systematic application to domain-specific knowledge graphs.
@@
Pre-trained models have been demonstrated to be effective in natural language processing, especially in data scarcity scenarios. Through training with limited data, pre-trained models can be transferred to a specific task while showing strong performance. In materials science, pre-trained model are also widely used. Early work, like Mat2Vec 361, learned word embeddings in materials literature without human supervision, showing a surprising capability to identify semantics in materials science text. This led to the development of material-specific language models such as MatSciBERT 146 and MaterialBERT 320. They collect a large corpus of materials science for continual training without changing the BERT architecture. Compared to vanilla BERT, they show superior performance in identifying material-related property data and phrases in the literature. This conclusion contributes to the emergence of various BERT variations that were particular to certain materials subfields, such as BatteryBERT 156 for battery-related material, OpticalBERT 426 for optical materials, PolyBERT 196 for polymers materials, and SteelBERT 354 for metallurgy materials. UniMat 263 propelled the field forward by merging multiple modes (atomic structure, XRD patterns, and composition) into strong joint embeddings. This made it easier to find novel materials. SmileyLlama 62 is further eolved with Direct Preference Optimization(DPO) to optimize the structure generation. Recent ideas deal with structural limitations: Huang et al. 154 proposed a self-supervised and multimodal pre-training technique that can predict performance based only on stoichiometry, without explicit understanding of the atomic structure. MELT189 was the first to employ curriculum-based continuous pre-training, which used semantic knowledge graphs to convert models from broad-domain knowledge to material-specific concepts in a systematic way.